# SpiroLLM: Finetuning Pretrained LLMs to Understand Spirogram Time Series with Clinical Validation in COPD Reporting

### v1

@@ METHODS
The application of artificial intelligence in the diagnosis of COPD has made significant progress in recent years. Early research primarily employed traditional machine learning methods, which relied on key metrics extracted from PFTs. To overcome this limitation, recent studies have shifted their focus to deep learning models capable of learning features directly from raw spirometry time-series data. For instance, in our previous work, we proposed the DeepSpiro model 5, which processes flow-volume sequences directly to detect COPD. Similarly, Bhattacharya et al. 8 utilized a Fully Convolutional Network (FCN) to analyze raw expiratory flow curves to distinguish between different structural phenotypes of COPD, achieving accuracy that significantly surpassed methods relying solely on traditional PFT metrics.
Another technical approach involves transforming the time-series waveforms into images for analysis by well-established Convolutional Neural Networks (CNNs). For example, the AI-PFT-Clin model developed by Eun-Tae Jeon et al. 9 improved the predictive accuracy for COPD acute exacerbation events by fusing clinical variables with images of flow-volume loops and volume-time curves. Eva Topole et al. 10 also converted flow-volume curves into low-resolution images and used a CNN to automatically assess the acceptability of spirometry maneuvers.
However, nearly all of these advanced research efforts adhere to a discriminative paradigm. Their core tasks are classification, prediction, or phenotyping. Their final outputs are discrete class labels or continuous risk scores, rather than coherent, narrative text that can explain the diagnostic rationale. This fundamentally limits the applicability of these models in clinical scenarios that require detailed diagnostic explanations.
@@
Concurrently, the emergence of Multimodal Large Language Models (MLLMs) has provided a powerful technological foundation for the automated generation of complex medical narrative texts 6. These models can integrate and comprehend medical data from diverse sources to produce high-quality reports. In their survey, Ye et al. 11 systematically summarized the applications of MLLMs in areas such as medical report generation, diagnosis, and treatment. These models are capable of processing and integrating data from multiple sources, including medical imaging, Electronic Health Records (EHRs), and laboratory results.
@@
In the field of medical imaging, MLLMs have been successfully applied to the automated generation of radiology reports. For instance, the MRG-LLM proposed by Li et al. 12 can generate more targeted and accurate reports for input X-ray images. Li et al. 13 later extended this technique to report generation for 3D brain CT scans. These studies demonstrate the robust capability of MLLMs to process complex visual information and translate it into specialized text. To address the challenge of scarce multimodal medical data, Chen et al. 14 proposed the AdaCoMed framework, which effectively enhances model performance through collaborative learning between large and small models. This work highlights a trend of evolving from simple feature concatenation towards deeper cross-modal interaction. Meanwhile, to address privacy concerns, Kaissis et al. 15 introduced a framework based on federated learning to enable multi-center model training while safeguarding data privacy.
Crucially, MLLMs have also shown significant potential in processing physiological time-series data, which is similar to the core data type in our study. The GEM model proposed by Lan et al. 16 successfully generated clinically interpretable ECG diagnostic reports through effective cross-modal fusion and alignment of electrocardiogram (ECG) time-series signals and images. These studies confirm the technical feasibility of generating complex diagnostic reports from multi-source, heterogeneous medical data, providing valuable architectural references for our research.
@@
In summary, despite the respective progress made in the discriminative analysis of pulmonary function data and in general-purpose medical report generation models, a significant gap exists between them. To date, no prior work has attempted to apply powerful generative models to the domain of pulmonary function diagnostics. Specifically, there is a lack of a framework capable of fusing the raw spirometry time-series with structured PFT metrics—two highly complementary modalities. Therefore, this study proposes SpiroLLM, which, by designing a novel fusion architecture, aims to extend the task of pulmonary function analysis for the first time from traditional classification and prediction to the more clinically valuable task of automated diagnostic report generation. This endeavor seeks to provide a new technical pathway for the fine-grained and interpretable diagnosis of COPD.
@@
To extract deep feature embeddings $E_{i}$ from the pulmonary function time-series $s_{i}$, this study adopts the DeepSpiro model proposed in prior work 5 as the core time-series encoder. This encoder, denoted as $E_{s}$ in the formula, employs a hybrid CNN-BiLSTM architecture. It first captures key local patterns in the sequence using a one-dimensional Convolutional Neural Network (1D-CNN) and subsequently models the temporal context of these local features using a Bidirectional Long Short-Term Memory (BiLSTM) network.
@@
[…]
@@
This step aims to obtain a qualitative description of the patient’s respiratory curve morphology. We first visualize the raw Flow-Volume time-series data to generate standard Flow-Volume curve images. Subsequently, we utilize a powerful multimodal large language model (Qwen2.5-VL-72B17), in conjunction with meticulously designed prompts aimed at guiding the model to focus on key morphological features of the curve (e.g., peak shape, degree of concavity in the expiratory limb, etc.), to automatically generate an accurate and objective textual description (see Appendix A for details). This description forms the foundation of the gold-standard report’s section on the patient’s individualized physiological presentation.
@@
To obtain quantitative clinical evidence, we developed a pulmonary function metric calculation tool named SpiroUtils. This tool directly processes the raw time-series data to precisely calculate a series of key pulmonary function parameters, including Forced Vital Capacity (FVC), Forced Expiratory Volume in the first second (FEV1), Forced Expiratory Flow between 25% and 75% of FVC (FEF25%-75%), and FEF75%. More importantly, SpiroUtils integrates the patient’s demographic information (age, gender, height) to calculate the Predicted Value and Z-score for these metrics based on the multi-ethnic reference equations published by the Global Lung Function Initiative in 201218.
@@
To ensure that the generated reports comply with the latest clinical guidelines, we constructed a domain knowledge base. The content of this knowledge base is derived from the GOLD 2025 Report1. During the report generation process, we employ Retrieval-Augmented Generation (RAG) techniques. Specifically, the morphological descriptions and quantitative metrics from the previous two steps are used as a composite query to retrieve the most relevant knowledge snippets—such as diagnostic criteria, severity grading, and treatment recommendations—from the knowledge base that correspond to the current patient’s condition.
@@
After obtaining the qualitative morphological descriptions of the respiratory curve, the precise quantitative physiological metrics, and the authoritative knowledge from the GOLD standards, we integrate these three components: the morphological description text, the quantitative metrics including Z-scores, and the retrieved relevant knowledge snippets. This integrated information is then formatted according to a meticulously designed structured prompt template (see Appendix B for details). Subsequently, this structured, comprehensive prompt is input into the DeepSeek-V3 model19 to finally generate a gold-standard diagnostic report $r_{i}^{*}$ that is comprehensive in content, reliable in its conclusions, and aligns with the linguistic style of clinicians. This high-quality report serves as the ground-truth label for the supervised fine-tuning of the main model, thereby ensuring that the model learns accurate diagnostic logic and professional expression.
Dataset
This study is based on a high-quality, large-scale cohort constructed through a multi-stage, rigorous screening process from the UKB, comprising a total of 234,028 individuals. During its construction, the cohort was subject to a stringent quality control process to ensure both physiological validity and high data precision. The key pulmonary function metrics for all selected samples—including Forced Expiratory Volume in 1 second (FEV1), Forced Vital Capacity (FVC), and Peak Expiratory Flow (PEF)—were required to meet pre-defined validity thresholds, and the relative error between their calculated values and the official UKB measurements was controlled to be within 10%. This screened cohort, which includes patients with definitive COPD diagnoses and non-COPD control individuals selected under the same strict criteria, provides an unprecedentedly large-scale data foundation for the training and validation of the SpiroLLM model in this study.
To further screen our high-quality, large-scale cohort of 234,028 individuals and select the "ground true" samples best suited for training the SpiroLLM model, we designed and implemented an innovative screening process. This process integrates generative artificial intelligence techniques with an automated quality assessment mechanism. Specifically, we utilized a large language model to generate preliminary diagnostic reports for the candidate samples and constructed an automated evaluation system to rigorously assess the quality of these reports in terms of informational completeness, logical coherence, and diagnostic accuracy. Based on this, only individual samples for which high-quality, high-fidelity diagnostic reports could be consistently generated were retained for inclusion in the final dataset used for model training. This screening mechanism effectively ensures the high purity and clinical consistency of the training data, providing high-quality supervision signals for the model’s learning process.
@@
For the purposes of model training, validation, and final performance evaluation, this high-quality, large-scale cohort was partitioned into training, validation, and test sets at an 8:1:1 ratio. To ensure that the ratio of cases to controls remained consistent across all data subsets, we employed stratified random sampling for this division.
Implementation Details
@@
This study adopts the parameter-efficient LoRA (Low-Rank Adaptation) strategy to fine-tune the model20. During training, the weights of both the SpiroEncoder and the LLM backbone are kept frozen, where only the parameters of the SpiroProjector and the LoRA adapters in the LLM are updated. The AdamW optimizer is used, with different learning rates set for the SpiroProjector and the LoRA modules. The overall learning rate schedule employs a cosine annealing strategy with a warm-up period, and the optimization objective is the standard language model loss function. To enhance efficiency, bfloat16 mixed-precision is utilized throughout the training process, and an early stopping mechanism is configured to prevent overfitting. The training was conducted on 4 NVIDIA RTX 4090 GPUs.
@@ Appendix D Required Fields for Label Extraction
Appendix E Expert Evaluation Results

### v2

@@ METHODS
The application of artificial intelligence in the diagnosis of COPD has made significant progress in recent years. As highlighted in recent reviews by Gompelmann et al. gompelmann2025ai, AI is increasingly transforming COPD management, evolving from simple phenotyping to complex decision support systems. Early research primarily employed traditional machine learning methods, which relied on key metrics extracted from PFTs.
To overcome this limitation, recent studies have shifted their focus to deep learning models capable of learning features directly from raw spirometry time-series data. For instance, in our previous work, we proposed the DeepSpiro model mei2025deep, which processes flow-volume sequences directly to detect COPD. More recently, Cosentino et al. cosentino2023inference validated this paradigm on a large scale using UK Biobank data, demonstrating that Convolutional Neural Networks (CNNs) trained on raw spirograms could identify novel genetic loci. Similarly, Hill et al. hill2025deep introduced Spiro-CLF, a contrastive learning framework that leverages even suboptimal spirometry maneuvers to predict mortality.
Another technical approach involves transforming the time-series waveforms into images for analysis. For example, the AI-PFT-Clin model developed by Jeon et al. jeon2025deep improved the predictive accuracy for COPD acute exacerbation events by fusing clinical variables with images of flow-volume loops. Topole et al. topole2023artificial also converted flow-volume curves into low-resolution images and used a CNN to automatically assess the acceptability of spirometry maneuvers.
However, nearly all of these advanced research efforts adhere to a discriminative paradigm. Their core tasks are limited to classification or risk scoring, lacking the capability to generate coherent, narrative text that can explain the diagnostic rationale.
@@
Concurrently, the emergence of Multimodal Large Language Models (MLLMs) has provided a powerful technological foundation for the automated generation of complex medical narrative texts liu2025application. These models can integrate and comprehend medical data from diverse sources to produce high-quality reports. In recent surveys, Ye et al. ye2025multimodal and Alsaad et al. alsaad2024multimodal systematically summarized the landscape of MLLMs in healthcare, highlighting their potential to integrate heterogeneous data sources—including Electronic Health Records (EHRs), imaging, and sensor data—to enhance diagnostic precision.
@@
In the field of medical imaging, MLLMs have already been successfully applied to the automated generation of radiology reports, serving as a precedent for other modalities. For instance, the MRG-LLM proposed by Li et al. li2025multimodal can generate targeted and accurate reports for input X-ray images, a technique later extended by Li et al. li2025towards to 3D brain CT scans. These studies demonstrate the robust capability of MLLMs to process complex visual information and translate it into specialized text. Additionally, advanced frameworks like AdaCoMed chen2025multi have explored deeper cross-modal interactions through collaborative learning, while Kaissis et al. kaissis2021end introduced federated learning approaches to safeguard data privacy during multi-center training.
Beyond imaging, significant strides have been made in applying MLLMs to physiological time-series data. Med-PaLM M tu2024towards demonstrated the feasibility of a generalist biomedical AI capable of handling diverse modalities, while HeLM (Belyaeva et al.) belyaeva2023multimodal established a benchmark for grounding LLMs in individual-specific health records.
Architecturally, our SpiroLLM shares structural parallels with these domain-specific MLLMs, employing a Connectivist paradigm—comprising a pre-trained modality encoder and an alignment projector—to bridge the gap between continuous biological signals and discrete textual reasoning. A pertinent example is the GEM model lan2025gemempoweringmllmgrounded, which similarly utilizes a specialized encoder to align ECG signals for clinical interpretation.
However, critical architectural distinctions exist between these diagnostic MLLMs and general time-series adapters. Frameworks like GPT4TS by Zhou et al. zhou2023one and Time-LLM jin2023time have successfully repurposed frozen LLMs for universal time-series analysis. Yet, these models typically utilize a patching mechanism to treat time-series segments as input tokens, optimizing the model primarily for forecasting or anomaly detection tasks. In contrast, they lack the domain-specific alignment required to interpret the global morphological semantics of clinical waveforms for diagnostic reasoning, a critical gap our SpiroLLM aims to fill.
@@
In conclusion, although advances have been made in both the discriminative analysis of pulmonary function data and the generation of general medical reports, a significant disconnect persists between these two domains. Specifically, powerful generative models have yet to be systematically applied to pulmonary functional diagnostics, and there is a notable absence of a unified framework capable of deeply integrating raw pulmonary function temporal waveforms with structured PFT metrics. To address this deficiency, we propose SpiroLLM. By designing a novel fusion architecture, this study advances the task of pulmonary function analysis from traditional classification prediction to the generation of automated, interpretable diagnostic reports. This work aims to establish a new technological pathway for the fine-grained, interpretable clinical diagnosis of COPD.
@@
To extract deep feature embeddings $E_{i}$ from the pulmonary function time-series $s_{i}$, this study adopts the DeepSpiro model proposed in prior work mei2025deep as the core time-series encoder. This encoder, denoted as $E_{s}$ in the formula, employs a hybrid CNN-BiLSTM architecture. It first captures key local patterns in the sequence using a one-dimensional Convolutional Neural Network (1D-CNN) and subsequently models the temporal context of these local features using a Bidirectional Long Short-Term Memory (BiLSTM) network.
@@
$\begin{split}\mathbf{P}_{i}&=\operatorname{SpiroProjector}\left(\mathbf{E}_{i}\mid\theta_{P}\right)=\operatorname{Dropout}\left(\operatorname{ReLU}\left(\mathbf{E}_{i}\mathbf{W}_{1}+\mathbf{b}_{1}\right)\right)\mathbf{W}_{2}+\mathbf{b}_{2}\end{split}$ (2)
@@
This step aims to obtain a qualitative description of the patient’s respiratory curve morphology. We first visualize the raw Flow-Volume time-series data to generate standard Flow-Volume curve images. Subsequently, we utilize a powerful multimodal large language model (Qwen2.5-VL-72Bbai2025qwen2), in conjunction with meticulously designed prompts aimed at guiding the model to focus on key morphological features of the curve (e.g., peak shape, degree of concavity in the expiratory limb, etc.), to automatically generate an accurate and objective textual description (see Appendix A for details). This description forms the foundation of the gold-standard report’s section on the patient’s individualized physiological presentation.
@@
To obtain quantitative clinical evidence, we developed a pulmonary function metric calculation tool named SpiroUtils. This tool directly processes the raw time-series data to precisely calculate a series of key pulmonary function parameters, including Forced Vital Capacity (FVC), Forced Expiratory Volume in the first second (FEV1), Forced Expiratory Flow between 25% and 75% of FVC (FEF25%-75%), and FEF75%. More importantly, SpiroUtils integrates the patient’s demographic information (age, gender, height) to calculate the Predicted Value and Z-score for these metrics based on the multi-ethnic reference equations published by the Global Lung Function Initiative in 2012quanjer2012multi.
@@
To ensure that the generated reports comply with the latest clinical guidelines, we constructed a domain knowledge base. The content of this knowledge base is derived from the GOLD 2025 ReportVenkatesan2025. During the report generation process, we employ Retrieval-Augmented Generation (RAG) techniques. Specifically, the morphological descriptions and quantitative metrics from the previous two steps are used as a composite query to retrieve the most relevant knowledge snippets—such as diagnostic criteria, severity grading, and treatment recommendations—from the knowledge base that correspond to the current patient’s condition.
@@
After obtaining the qualitative morphological descriptions of the respiratory curve, the precise quantitative physiological metrics, and the authoritative knowledge from the GOLD standards, we integrate these three components: the morphological description text, the quantitative metrics including Z-scores, and the retrieved relevant knowledge snippets. This integrated information is then formatted according to a meticulously designed structured prompt template (see Appendix B for details). Subsequently, this structured, comprehensive prompt is input into the DeepSeek-V3 modelliu2024deepseek to finally generate a gold-standard diagnostic report $r_{i}^{*}$ that is comprehensive in content, reliable in its conclusions, and aligns with the linguistic style of clinicians. This high-quality report serves as the ground-truth label for the supervised fine-tuning of the main model, thereby ensuring that the model learns accurate diagnostic logic and professional expression.
Dataset Construction
This study leverages data from 453,558 participants in the UK Biobank (UKB). To ensure the physiological validity and reliability of subsequent analyses and to minimize potential confounding factors, a systematic quality control and sample screening process was implemented. The detailed screening workflow is presented in Appendix Figure E.
Initially, Chronic Obstructive Pulmonary Disease (COPD) cases were defined based on clinical diagnostic criteria (ICD-10 codes J43, J44, and self-reported illness codes, as detailed in Appendix D) rather than a single spirometric threshold, ensuring the definition comprehensively reflects clinical reality. Concurrently, rigorous data quality controls were applied: participants of non-European descent were excluded to minimize population stratification; records with invalid spirometry tests were removed; statistical outliers in key pulmonary function metrics (top/bottom 0.5%) were eliminated; and unreliable data were excluded based on signal inconsistency between repeated measurements (relative difference > 10%). Through this parallel screening and definition process, a baseline valid cohort of 234,028 individuals was obtained, comprising 10,748 confirmed COPD cases, with a case-to-control ratio of approximately 1:20.
It is recognized that relying on clinical diagnostic codes may introduce label uncertainties, and the baseline cohort exhibited significant class imbalance. To address these issues, a balanced subset was first constructed by retaining all COPD cases and randomly downsampling healthy controls, resulting in a balanced dataset of 21,496 samples (1:1 ratio). Subsequently, to enhance label reliability, an innovative AI-assisted quality assurance mechanism was introduced. This involved utilizing a Large Language Model to perform a consistency check between the clinical diagnosis and physiological evidence (spirometry results) for each sample. This step excluded 3,080 samples where significant contradictions were identified, thereby effectively mitigating potential label inconsistencies and ensuring rigorous alignment between clinical diagnoses and physiological data.
@@
The final refined dataset comprises 18,416 high-quality samples, consisting of 8,245 COPD cases and 10,171 healthy controls. Using stratified random sampling, the dataset was partitioned into training ($N=14,730$), validation ($N=1,843$), and test ($N=1,843$) sets at an 8:1:1 ratio. This partitioning strategy ensured consistent class distribution across all subsets, providing a reliable foundation for subsequent model training and evaluation.
Experimental Details
@@
This study adopts the parameter-efficient LoRA (Low-Rank Adaptation) strategy to fine-tune the modelhu2022lora. During training, the weights of both the SpiroEncoder and the LLM backbone are kept frozen, where only the parameters of the SpiroProjector and the LoRA adapters in the LLM are updated. The AdamW optimizer is used, with different learning rates set for the SpiroProjector and the LoRA modules. The overall learning rate schedule employs a cosine annealing strategy with a warm-up period, and the optimization objective is the standard language model loss function. To enhance efficiency, bfloat16 mixed-precision is utilized throughout the training process, and an early stopping mechanism is configured to prevent overfitting. The training was conducted on 4 NVIDIA RTX 4090 GPUs.
Baseline Models
To rigorously evaluate the proposed method, a comprehensive benchmarking framework was established, categorized into three distinct groups.
Demographic and Anthropometric Baselines were constructed using XGBoost in a hierarchical manner: XGBoost-Demo utilized only basic demographics (Age, Sex, and Smoking Status); XGBoost-Anthro incorporated anthropometric data (Height and BMI) into the demographic baseline; and XGBoost-Clinical further added quantitative spirometry metrics derived from SpiroUtils (including FEV1, FVC, FEV1/FVC ratio, PEF, and FEF25-75), establishing a robust benchmark reflecting comprehensive clinical assessment.
Time-Series Baselines compared the proposed approach against state-of-the-art waveform encoders, including the flow-volume ResNet18 proposed by Cosentino et alcosentino2023inference., advanced time-series models (Informerzhou2021informer, PatchTSTYuqietal-2023-PatchTST, and TimesNetwu2022timesnet), and SpiroEncoder, the feature extraction backbone employed in the previously developed DeepSpiro frameworkmei2025deep. These models rely exclusively on raw flow-volume curve data.
The Multimodal Fusion Baseline featured DeepSpiromei2025deep, which combines raw flow-volume curves with basic demographics and the FEV1/FVC ratio. This model serves as a representative baseline for traditional feature concatenation strategies, facilitating a rigorous evaluation of the performance gains achieved by the proposed method.
@@ Appendix D Required Fields for Label Extraction
Appendix E Data Preprocessing Flowchart
Figure 5: Data Preprocessing Pipeline.
Appendix F Expert Evaluation Results